OpenStudy (anonymous):

The function f(x) = 582(7)x represents the growth of a mosquito population every year in a remote swamp. Troy wants to manipulate the formula to an equivalent form that calculates every 2 months, not every year. Which function is correct for Troy's purposes?

OpenStudy (solomonzelman):

Got disconnected, apologize. Your monthly rate is going to be 12 times smaller than the yearly rate. \(\large\displaystyle f(x) = 582(7)^{ x }\) is a yearly growth. \(\large\displaystyle f(x) = 582(7)^{ \left(^x/_{12}\right) }\) is the monthly rate.

OpenStudy (solomonzelman):

same way, you can rewrite it as: \(\large\displaystyle f(x) = 582(\sqrt[12]{ 7})^{ x}\)

OpenStudy (anonymous):

but every 2 months